OpenStudy (anonymous):

Factor completely: 5x^2 - 40 would it be 5(x + 8)(x - 8)?

OpenStudy (anonymous):

Right :)

OpenStudy (anonymous):

You use first Common monomial then Factor the remaining terms

OpenStudy (anonymous):

It would be 5(x + 2)(x - 2)

OpenStudy (ybarrap):

Whenever you have something like \( a^{2} -b^{2}\) then you can factor it like \( (a+b)(a-b)\) since you have \(5x^{2}-40\) we can rewrite as \(5(x^2-8)\). This is now in the form we can use the technique I just described.: \(5(x^{2}-\sqrt{8}^{2})=5(x+\sqrt{8})(x-\sqrt{8})=5(x+2\sqrt{8})(x-2\sqrt{8})\).